Briefly explain exchange rate theories: Interest Rate Parity (IRP) and Purchasing Power Parity (PPP) and the International Fisher Effect (IFE). How do these work?
Answer #1

A) Interest rate parity theory states that the difference between the interest rate of two countries is equal to the difference calculated by forward and spot exchange rate . This is a no arbitrage condition, where the investor will be indifferent in borrowing from banks in either country.

Formula : F0 = S0 ( 1+ Ra) / (1+Rb)

F0 = forward rate. , S0 = Spot rate , Ra = interest rate in country A. , Rb = interest rate in country B.

B) Purchasing Power Parity is a theory which is based on law of one price. According to this theory a good should be worth same in both the countries , taking into account the exchange rate. If this does not happen then there will be imbalance between the countries , and the country having the advantage will exploit the other. It is applicable only in long run.

C) International fisher effect is an extention of fisher effect and is used in foreign exchange trading . This theory states that the difference between the the exchange rate of two currencies is equal to the difference between the nominal interest rate of the individual countries.

According to this theory if a country has higher interet rate , then it will experience higher inflation as compared to the country with lower interest rate.

Formula : E = R1 - R2 / 1+ R2 = R1 - R2

E = percent change in exchange rate

R1 = nominal interest rate of country 1

R2 = nominal interest rate of country 2